OpenStudy (anonymous):

how do i graph systems of equations by graphing for example x-y=2 x+y=6

OpenStudy (anonymous):

Get these two equations in term of y

OpenStudy (anonymous):

y=x-2 y=-x+6

OpenStudy (anonymous):

Now plot two line above and see where they meet

myininaya (myininaya):

2x=8 x=4 4-y=2 -y=-2 y=2

myininaya (myininaya):

graph he lines imranmeah has then find where they meet (or intersect)

myininaya (myininaya):

and you should get they meet at (4,2)